OpenStudy (anonymous):

Evaluate the radical expressions or indicate that the root is not a real number. ^3 sq root (-4)^3

Miracrown (miracrown):

\[\sqrt[3]{(-4)^{3}}\] So it is a cube root? That's the part I'm unsure about I assume ^3 sq root means cube root, but you never know

OpenStudy (anonymous):

Yes

Miracrown (miracrown):

Okie... well, remember what a cube root is: In this case, they want: "The number that, when we raise it to the 3rd power, we get (-4)^3" So what do we have to raise the 3rd power, to get (-4)^3?

Miracrown (miracrown):

\[?^{3} \space = (-4)^3\]

OpenStudy (anonymous):

I'm not sure! Would it be -2?

Miracrown (miracrown):

It would actually be -4 You can think of it as the 3rd root and 3rd power (inside the root) "cancelling" each other

OpenStudy (anonymous):

Oh ok, I understand

OpenStudy (anonymous):

So would that be my answer?

Miracrown (miracrown):

\[(-4)^3 = (-4)^3\] Yep, that's what it boils down to! So the root itself is -4

Miracrown (miracrown):

-4 is the answer here: It's the cube root of (-4)^3

OpenStudy (anonymous):

Thank you so much!

Miracrown (miracrown):

You're welcum =]

OpenStudy (anonymous):

Can you help me with this one too? Solve the linear equation. 2(x + 6) + 6 = 3(x + 5) + 7 A. {6} B. {-4} C. {12} D. {9}

Miracrown (miracrown):

Sure. What do we get if you distribute the 2's and the 3?